OpenStudy (anonymous):

Given f(x) = (3x-1)/2 Solve for f^1 (4)

OpenStudy (anonymous):

@jim_thompson5910 @Hero @phi @zepdrix @.Sam. @timo86m @jdoe0001

zepdrix (zepdrix):

\[\Large f(x)=\frac{3x-1}{2}\] Solve for `f^1(4)` Did you mean to write \(\Large f^{-1}(4)\) ?

OpenStudy (anonymous):

No, its: \[f ^{1} (4)\]

zepdrix (zepdrix):

What is the 1 suppose to signify? Is that a first derivative? :o

OpenStudy (anonymous):

I have no idea, that was the question I was given.

zepdrix (zepdrix):

Hmm, weird. :(

OpenStudy (anonymous):

The answer choices are: a. 1 b. 3 c. 5 d. 7

OpenStudy (ness9630):

What class are you in?

OpenStudy (anonymous):

Algebra 2

OpenStudy (ness9630):

Well it can't mean derivative, that's for sure.

OpenStudy (anonymous):

So how do I solve this then?

OpenStudy (ness9630):

@zepdrix I know you're smart.. any ideas?

zepdrix (zepdrix):

I think it was a typo, I think it was supposed to be f^-1(-4). Was this from a book, or like something your teacher assigned? :o Have you learned about finding inverses yet troy?

OpenStudy (ness9630):

Wait.. hold on! I think it was suppose to be f^-1 like zep says.. it gets you one of the answer choices you listed. Watch, try it: \[f(x)=\frac{3x-1}{2}\] \[x=\frac{3y-1}{2}\] Solve for y and you get this: \[2x=3y-1\] \[2x+1=3y\] Divide: \[\frac{2x+1}{3}=f^{-1}(4)\] Can you finish it?

OpenStudy (anonymous):

I like just started this course, I haven't really learned much yet. It's mostly just review of Algebra 1 from a couple years ago, which I can't remember anyway.

OpenStudy (ness9630):

Do you understand what I just did though?

OpenStudy (anonymous):

No.. not really.

OpenStudy (ness9630):

What I did first was switch y and x around, since that's what solving for an inverse is. Then I just solved for y again.

OpenStudy (anonymous):

Ok, so how would I get to the answer?

OpenStudy (ness9630):

You see that very last equation? f^-1(4) Just plug in 4 for x and you will get it.

OpenStudy (anonymous):

So, 3?

OpenStudy (ness9630):

Yup!

OpenStudy (anonymous):

Thanks! I've got another for you, if you don't mind. On the graph of the equation 4x + 5y = 2, what is the value of the y-intercept?

OpenStudy (ness9630):

The y-intercept is just when x=0

OpenStudy (anonymous):

The choices are: a. -5 b. -4 c. 4 d. 5

OpenStudy (ness9630):

That's strange.. it's not one of the choices

OpenStudy (anonymous):

The answer is 4, you have to graph the equation and tell what number is on the y-intercept.

OpenStudy (anonymous):

The next question is: Graph: tinyurl.com/mrr8dcb Find the average rate of change for the given function from x=1 to x=2.

OpenStudy (ness9630):

I have no access to connexus

OpenStudy (anonymous):

Sorry, one second

OpenStudy (anonymous):

here:

OpenStudy (ness9630):

Just find the slope: (1, 3) (2, 6)

OpenStudy (anonymous):

4 isn't and option.

OpenStudy (anonymous):

a. -2 b. -1/2 c. 1/2 d. 2

OpenStudy (ness9630):

Whoops, it's suppose to be (1, 3)(2, 5)

OpenStudy (anonymous):

Oh, ok

OpenStudy (anonymous):

So it's 2?

OpenStudy (ness9630):

Yup! Unless I'm losing my skills

OpenStudy (anonymous):

haha hopefully not! Ok, I got another one. (Use the table below to answer the question) Find the average rate of change for the given function from x=1 to x=3

OpenStudy (ness9630):

I think for that one you would find the slopes for each, then add and divide the slopes

OpenStudy (anonymous):

Ok. Thanks for all your help!

OpenStudy (ness9630):

Anytime!